Design a circuit that adds together 4X and 2X to give 6X, where X is X2X1X0, and gives a 6-bit sum S5S4S3S2S1S0. This can be done using 2 half adders and 1 full adder, no additional gates are required.
Answer #1

Q We have to add 4X+ 2X = 6X Here x 44, 14Yo G bit oul put is S5S4835,5,5o. Gbit Equivalent of x = 000X2X1 Xo 6 bit EquivalenSo our enpressions are given below l- here require two haff odders and 1 full adder. So=0 Si=xo S2 = Sum (X1, Xo) Sz= Sum (X2